Let x and y be real numbers. If (x+yi) - (2-3i) = -6+4i, what are the values of x and y?
a) x=8, y=7
b) x=8, y=1
c) x=-4, y=7
d) x=-4, y=1

Answer by sudhanshu_kmr(1152)


(x + yi)- (2-3i) = -6 + 4i
=> x-2 + i(y+3) = -6 + 4i

by comparing real and imaginary part ........
x = -6 + 2 = -4
y = 4 -3 = 1

(d) is correct ......